OpenStudy (simplymarie_x):

How many liters of water vapor can be produced if 8.9 liters of methane gas (CH4) are combusted, if all measurements are taken at the same temperature and pressure? Show all of the work used to solve this problem. CH4 (g) + 2 O2 (g) ---> CO2 (g) + 2 H2O (g)

OpenStudy (anonymous):

Hint: What is the moles of methane gas combusted?

OpenStudy (photon336):

are you given any other information?

OpenStudy (anonymous):

The information is sufficient to solve the problem

OpenStudy (simplymarie_x):

I have no idea how to do this at all. But that's all that is given to me. @Photon336 @thushananth01

OpenStudy (anonymous):

Moles of a gas = Volume of gas/ molar volume = 8.9 /24

OpenStudy (anonymous):

The ratio of moles between CH4 and H20 = 1 :2

OpenStudy (photon336):

yeah, this could be based off the fact that 1 mole of every gas occupies 22.4L at STP, then you could easily find out the number of moles from this.

OpenStudy (anonymous):

@Photon336 That's right

OpenStudy (photon336):

@simplymarie_x follow what @thushananth01 said and then tell us what you get.

OpenStudy (simplymarie_x):

So 8.9/24 = 0.397 moles then since the ratio is 1:2, I'd do 0.397*2 = 0.794 then I'd have to convert it back to liters so 0.794 * 22.414 giving me an answer of 17.8 L of water vapor?

OpenStudy (anonymous):

Divide 8.9 by 22.4 and then use that moles to find moles of water vapor

OpenStudy (simplymarie_x):

So I'm not right? Couldn't you also just do the 8.9 liters of methane gas x 2 to get the answer?

OpenStudy (simplymarie_x):

which would be 17.8 ? @thushananth01 @photon336

OpenStudy (anonymous):

That's right! Good job!

OpenStudy (simplymarie_x):

Okay thank you so much! :D

OpenStudy (anonymous):

Assuming that it is at standard room temperate and pressure